OpenStudy (anonymous):

3. Define a transformation from P2 to R3 by: \[p(x) -> ( \int\limits_{0}^{1} p(x) d(x) , \int\limits_{0}^{2} p(x) d(x) , \int\limits_{0}^{3} p(x) d(x) )\] (a) Is this a homomorphism? (b) Is it onto? (c) Is it one-to-one? (d) Is it an isomorphism? (e) Find a matrix for the transformation between the representation of the standard basis of P2 and the standard basis of R3.

OpenStudy (anonymous):

2 spaces

OpenStudy (anonymous):

Basically transforming a vector with two entries to a vector with 3 entries right?

OpenStudy (anonymous):

This is a linear transformation right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

do you have an idea?!

OpenStudy (anonymous):

help @eliassaab

OpenStudy (anonymous):

It is linear since integration is linear.

OpenStudy (anonymous):

so what 's the answer for these questions?

OpenStudy (anonymous):

Well the least you can do is write down the definition for each one. If no one answers then just ask your prof.

OpenStudy (anonymous):

It is onto since if you take \[ (u,v,w)\in R^3\\ p(x)=3 u - (3 v)/2 + w/3 + (-5 u + 4 v - w) x + 1/2 (3 u - 3 v + w) x^2 \] then the image of p(x) is (u,v,w)

OpenStudy (anonymous):

It is one to one since both spaces are of dimension 3 and there is an onto map from one to the othr

OpenStudy (anonymous):

I thought you couldnt have both.

OpenStudy (anonymous):

Thank you prof. @eliassaab

OpenStudy (anonymous):

The image of the polynomial 1 is {1,2,3} The image of the polynomial x is\[ \left\{\frac{1}{2},2,\frac{9}{ 2}\right\} \] The image of the polynomial x^2 is\[ \left\{\frac{1}{3},\frac{8}{3} ,9\right\} \]

OpenStudy (anonymous):

The matrix of the transformation is \[ \left( \begin{array}{ccc} 1 & \frac{1}{2} & \frac{1}{3} \\ 2 & 2 & \frac{8}{3} \\ 3 & \frac{9}{2} & 9 \\ \end{array} \right) \]